OpenStudy (anonymous):

It is given that (x+y)^2=16, (y+z)^2=36 and (z+x)^2=81 If x+y+z>3, then the number of possible value for x+y+z is

OpenStudy (asnaseer):

(x+y)^2=16 => x+y = +/- 4 - (a) (y+z)^2=36 => y+z = +/- 6 - (b) (z+x)^2=81 => z+x = +/- 9 - (c) (a)+(b): x+2y+z = +/-4 +/-6 - (d) (a)+(c): 2x+y+z = +/-4 +/-9 - (e) (b)+(c): x+y+2z = +/-6 +/-9 - (f) (d)+(e)+(f): 3x+3y+3z = 2(+/-4 +/-6 +/-9) therefore: x + y + z = 2(+/-4 +/-6 +/-9)/3 that should help you arrive at the answer.

OpenStudy (asnaseer):

sorry - I think I made a mistake there, you can just add (a)+(b)+(c) to get: 2x + 2y + 2z = +/-4 +/-6 +/- 9 x+y+z = (+/-4 +/-6 +/-9)/2

OpenStudy (asnaseer):

then throw away all results that do not satisfy: x+y+z > 3

OpenStudy (anonymous):

i don't understand, the way u write it down :(

OpenStudy (asnaseer):

\[x+y+z=\frac{\pm4\pm6\pm9}{2}\]

OpenStudy (asnaseer):

does that help?

OpenStudy (anonymous):

A Mathematica solution is attached.

OpenStudy (asnaseer):

@pinky_cute1995 - which part is it that you do not understand?

OpenStudy (anonymous):

the last part : (d)+(e)+(f): 3x+3y+3z = 2(+/-4 +/-6 +/-9) therefore: x + y + z = 2(+/-4 +/-6 +/-9)/3

OpenStudy (asnaseer):

ignore that - I made a mistake there. If you just add equations (a), (b) and (c) then you get:\[2x + 2y + 2z = \pm4 \pm6 \pm9\]therefore:\[2(x+y+z)=\pm4 \pm6 \pm9\]

OpenStudy (asnaseer):

do you understand this step:\[(x+y)^2=16\implies x+y=\pm4\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so if we do this for all three of your equations we end up with:\[(x+y)^2=16\implies x+y=\pm4\tag{a}\]\[(y+z)^2=36\implies y+z=\pm6\tag{b}\]\[(z+x)^2=81\implies z+x=\pm9\tag{c}\]understand so far?

OpenStudy (anonymous):

yah

OpenStudy (asnaseer):

ok, so now we just add all equations (a)+(b)+(c) to get:\[x+y+y+z+z+x=\pm4\pm6\pm9\]therefore:\[2x+2y+2z=\pm4\pm6\pm9\]therefore:\[2(x+y+z)=\pm4\pm6\pm9\]make sense?

OpenStudy (anonymous):

oh

OpenStudy (asnaseer):

so now make use of the fact you are given:\[x+y+z>3\]therefore:\[2(x+y+z)>6\]so see which of the combinations of:\[\pm4\pm6\pm9\]give you a value >6